A study looking at ovarian cancer in women compared cases with non-cases, and found that 150/200 cases did not use birth control compared with 50/200 of the non-cases.

1) Create a table to display the data.

2) Calculate the odds of exposure in cases and non-cases.

3) Calculate the odds ratio using the cross-product ratio.

4) How does the difference between the two prevalences of ovarian cancer (75% vs 25%) compare to the odds ratio?
Answer #1

Answer 1 :)

Subjects

Exposed

Unexposed

Total

Cases

150

50

200

Non Cases

50

150

200

Total

200

200

400

Answer 2 :)

Odds of Exposure in the case= Exposed/Unexposed

Odds of Exposure in the case = 150/50

Odds of Exposure in case = 3

Odds of Exposure in the non-case= Exposed/Unexposed

Odds of Exposure in the non-case = 50/150

Odds of Exposure in the non-case = 0.3

Answer 3 :)

Odds ratio = Odds of Exposure in the case /Odds of Exposure in the non-case

Odds ratio = 3 /0.3

Odds ratio = 10:1

Answer 4:)

The ratio of the two prevalence of ovarian cancer = 75% / 25%

The ratio of the two prevalence of ovarian cancer = 3

The prevalence ratio is 3:1 while the odds ratio is 10:1. It suggests that the prevalence does not provide sufficient evidence of the severity of the disease while the odds ratio is suggesting the risk factor associated with the cases.

""""""""""Please rate the answer"""""""""""